# Stability of a passive viscous droplet in a confined active nematic liquid crystal

### v1

@@ III.6.2 Deformation modes $(m\geq 2)$
$\begin{split}\Xi^{e}_{m}(\eta,\lambda)&=m\left[(m-1)\eta^{m-4}\,\overline{A}_{m}+(m+1)\eta^{m-2}\,\overline{C}_{m}\right]\\[5.0pt] &\,\,\,\,\,\,-\frac{m^{2}\eta^{m-2}}{\eta^{2m}-1}\left\{(m-1)^{2}\left[\frac{1-\eta^{2m-2}}{m-1}+1-\eta^{-2}\right]A_{m}-(m+1)^{2}\left[\frac{1-\eta^{-2m-2}}{m+1}-1+\eta^{-2}\right]B_{m}\right.\\[1.0pt] &\,\,\,\,\,\,\left.+(m^{2}-1)\left[\frac{1-\eta^{2m}}{m}+2\ln\eta\right]C_{m}-(m^{2}-1)\left[\frac{1-\eta^{-2m}}{m}-2\ln\eta\right]D_{m}\right\},\end{split}$ (63)
@@
Figure 5: (a) Color map of the total growth rate $\Lambda_{2}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=0.01$ and Ca $=100$. Linearly stable and unstable regions are shown in blue and red, respectively, and the grey curve shows the marginal stability curve. Effect of (b) $m$ and (c) Ca on the marginal stability curve in (a). Panels (d), (e), and (f) show the same as (a), (b), and (c) for a contractile system ($S=-1$).
@@
The dependence of $\Xi^{c}_{m}$, $\Xi^{a}_{m}$ and $\Xi^{e}_{m}$ on $m$, $\eta$ and $\lambda$ is illustrated in Fig. 4, where panels (a), (b) and (c) show the dependence on drop size $\eta$ for distinct modes $m$, and panels (d), (e) and (f) show the effect of viscosity ratio $\lambda$ for mode $m=2$. Capillary stresses always favor a circular drop and are therefore stabilizing, especially in the case of small drops and high wavenumbers; this damping effect of surface tension is most pronounced in low-viscosity drops. In the case of extensile systems, activity is destabilizing in the case of drops of small to intermediate sizes ($0\leq\eta\lesssim 0.5$), and stabilizing for large drops, and the range of unstable sizes increases with $m$; these trends are reversed for contractile systems. In both cases, low-viscosity drops are the ones that are most affected by activity (largest magnitude of the growth rate), and the limit of high-viscosity drops ($\lambda\to\infty$) is neutrally stable. On the other hand, nematic elasticity always stabilizes small to medium-sized drops, but weakly destabilizes large drops. These effects are most pronounced at low wavenumbers and low viscosity ratios.
@@
Figure 5(a) shows a color plot of the total growth rate $\Lambda_{2}$ corresponding to mode $m=2$ in the $(\eta,\lambda)$ parameter space, for an extensile system. The system is stable at lower values of the radius ratio ($0\leq\eta\leq 0.1$) followed by a range of unstable droplet sizes $(0.1<\eta\leq 0.25)$, and eventually regains stability at larger values of $\eta$. This behavior can be explained as follows: at lower values of $\eta$, capillary and elastic stresses dominate and are stabilizing; as $\eta$ increases, their effect weakens and the system becomes unstable as a result of active stresses; at larger values of $\eta$, active stresses become stabilizing leading to a negative growth rate. The marginal stability curve demarcating these regimes is only weakly affected by the viscosity ratio. As shown in Fig. 5(b), increasing the mode number $m$ tends to increase the size of the unstable region. Expectedly, decreasing the capillary number tends to stabilize the system across all drop sizes, especially at low viscosity ratios as illustrated in Fig. 5(c). The contractile case is analyzed similarly in Fig. 5(d,e,f). In this case, alternating regions of stable and unstable growth rates are not observed; rather, a single stable region at low values of $\eta$ gives way to an unstable region in the case of larger drops.
@@
$f^{e}_{rm}=-\frac{m^{2}\eta^{m-1}}{\eta^{2m}-1}\bigg{(}(2m-1)r^{m-3}+\frac{2m+1}{r^{m+3}}\bigg{)}\cos m\theta,\qquad f^{e}_{\theta m}=-\frac{m^{2}\eta^{m-1}}{\eta^{2m}-1}\bigg{(}r^{m-3}+\frac{1}{r^{m+3}}\bigg{)}\sin m\theta.$ (65)
They are plotted in Fig. 6 for mode $m=5$ and $\eta=0.45$, along with the first-order surface tractions $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a,e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a,e}$ acting on the interface. The fields shown in the figure suggest a destabilizing effect of bulk stresses and a stabilizing effect of surface tractions, both elastic and active (in the extensile case). The interplay of these forces, together with hydrodynamic stresses whose distribution is not explicitly known, gives rise to the trends found in Figs. 4 and 5.
Figure 6: (a) Active and (b) elastic bulk force densities and surface tractions at linear order in $\epsilon$ for the deformation mode $m=5$, in the case of a drop with radius $\eta=0.45$. In each case, the left plot shows the first-order force density, $\mathbf{f}^{a,e}_{1}=\nabla\cdot\bm{\sigma}^{a,e}_{1}$, in the bulk of the nematic, while the right plot shows the first-order traction field, $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a,e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a,e}$, on the drop surface. Panel (a) shows the active force and traction for an extensile system with $S=+1$.
@@ IV.1 Governing equations, boundary conditions and non-dimensionalization
$\mathbf{n}\cdot(\bm{\sigma}^{h}-\overline{\bm{\sigma}}^{a}-\overline{\bm{\sigma}}^{e}-\overline{\bm{\sigma}}^{h})=\frac{1}{\mathrm{Ca}}\kappa\mathbf{n}.$ (66)
@@ IV.2.1 Base state
In the base state, the nematic field lines inside the drop are also circles centered at the origin, with $\beta_{0}$ given by Eq. (18). A topological defect of charge $+1$ is present at the origin. The base state active and elastic stresses inside the drop are still given by Eq. (19). They induce a pressure field inside the drop that diverges at the location of the defect, while the pressure outside the drop is uniform:
$\overline{p}_{0}(r)=S\ln r-\frac{\xi^{2}}{2r^{2}},\qquad p_{0}=S\ln\eta+\frac{\xi^{2}}{2\eta^{2}}-\frac{1}{\eta\mathrm{Ca}}.$ (67)
@@ IV.2.2 Domain perturbation and linearized problem
$\beta_{1}(r,\theta)=\widetilde{\zeta}_{1}m\eta^{m-1}r^{m}\sin m\theta.$ (68)
@@
$\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}-\overline{\bm{\sigma}}^{a}_{1}-\overline{\bm{\sigma}}^{e}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}=\frac{\kappa_{1}\hat{\mathbf{r}}}{\mathrm{Ca}}-\frac{\widetilde{\zeta_{1}}}{\eta}\left(S-\frac{\xi^{2}}{\eta^{2}}\right)\big{(}\cos m\theta\,\hat{\mathbf{r}}+m\sin m\theta\,\hat{\bm{\theta}}\big{)}\qquad\mathrm{at}\,\,\,r=\eta.$ (69)
@@ IV.2.3 Lorentz reciprocal theorem
Application of the Lorentz reciprocal theorem follows the same step as discussed in Sec. III.5 and uses the same auxiliary Newtonian flow problem. Once again, the main difference with Sec. III.5 is that active and elastic stresses now appear inside the drop. As a result, the analog of Eq. (43) has no bulk contribution and reads
@@
$\int_{\mathcal{C}_{0}}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}-\overline{\bm{\sigma}}^{e}_{1}-\overline{\bm{\sigma}}^{a}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}\cdot\mathbf{U}\,\mathrm{d}\ell=m\pi\Sigma_{0}\widetilde{\zeta}_{1}\left[\frac{m^{2}-1}{\mathrm{Ca}}\frac{\overline{G}_{m}(\eta)}{\eta^{2}}-\left(S-\frac{\xi^{2}}{\eta^{2}}\right)\left(\frac{\overline{G}_{m}(\eta)}{\eta}-\overline{G}_{m}^{\prime}(\eta)\right)\right].$ (73)
@@
$\begin{split}\int_{\Omega_{0}^{-}}\big{(}\overline{\bm{\sigma}}^{e}_{1}+\overline{\bm{\sigma}}^{a}_{1}\big{)}:\overline{\mathbf{E}}\,\mathrm{d}S&=-\pi\Sigma_{0}\widetilde{\zeta}_{1}m\eta^{m-1}\left\{S\int_{0}^{\eta}r^{m}\bigg{[}r\overline{G}_{m}^{\prime\prime}(r)-\overline{G}_{m}^{\prime}(r)+m^{2}\frac{\overline{G}_{m}(r)}{r}\bigg{]}\mathrm{d}r\right.\\ &\,\,\,\,\,\,-\xi^{2}\int_{0}^{\eta}r^{m}\bigg{[}\frac{\overline{G}_{m}^{\prime\prime}(r)}{r}-(2m^{2}+1)\frac{\overline{G}_{m}^{\prime}(r)}{r^{2}}+3m^{2}\frac{\overline{G}_{m}(r)}{r^{3}}\bigg{]}\mathrm{d}r\bigg{\}}.\end{split}$ (74)
@@ IV.3 Dispersion relation and growth rates
Inserting Eqs. (46), (73) and (74) into Eq. (72) provides an expression for the growth in the form on Eq. (52), with contributions from capillary, active and elastic stresses that we discuss next.
In the translational mode ($m=1$), capillary stresses do not play a role and the active and elastic growth rates are
$\Xi_{1}^{a}(\eta,\lambda)=(2\eta-\eta^{3})\overline{A}_{1},\qquad\Xi_{1}^{e}(\eta,\lambda)=-\frac{2}{\eta}\overline{A}_{1}.$ (75)
These growth rates are plotted in Fig. 7 vs $\eta$ and $\lambda$. Active stresses are destabilizing in extensile systems and stabilizing in contractile systems, whereas elasticity always resists translation. Both active and elastic effects are most pronounced for small drops and at low viscosity ratios. In extensile systems with finite elasticity, the stability of the system in the $(\eta,\lambda)$ parameter space is shown in Fig. 7(c): small drops are all stabilized by elasticity, whereas drops with intermediate size can be unstable under translation as a result of activity.
@@
$\Xi_{m}^{a}(\eta,\lambda)$ $=m\left[(m-1)\eta^{3m-2}\overline{A}_{m}+m\eta^{3m}\overline{C}_{m}\right]+m\left[(m-1)\eta^{m-2}\overline{A}_{m}+(m+1)\eta^{m}\overline{C}_{m}\right],$ (77)
$\Xi_{m}^{e}(\eta,\lambda)$ $=m(m-1)\left[m\eta^{3m-4}\overline{A}_{m}+(m+1)\eta^{3m-2}\overline{C}_{m}\right]-m\left[(m-1)\eta^{m-4}\overline{A}_{m}+(m+1)\eta^{m-2}\overline{C}_{m}\right].$ (78)
@@
Figure 7: (a) Active contribution $\Xi^{a}_{1}$ and (b) elastic contribution $\Xi^{e}_{1}$ to the growth rate $\Lambda_{1}$ for the translational mode ($m=1$), plotted as functions of $\eta$ for different viscosity ratios $\lambda$. (c) Color map of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=1$. Linearly stable and unstable regions are shown in blue and red, respectively, and the grey curve shows the marginal stability curve.
[…]

### v2

@@ III.6.2 Deformation modes $(m\geq 2)$
$\begin{split}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi^{e}_{m}(\eta,\lambda)}&={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m\left[(m-1)\bigg{(}\frac{m^{2}(\eta^{2m}+1)}{\eta^{2m}-1}+m+1\bigg{)}\eta^{m-4}\,\overline{A}_{m}+\bigg{(}\frac{m^{3}(\eta^{2m}+1)}{\eta^{2m}-1}-2m-1\bigg{)}\eta^{m-2}\,\overline{C}_{m}\right]}\\[5.0pt] &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\,\,\,\,\,+\frac{2m^{2}\eta^{m-2}}{\eta^{2m}-1}\bigg{[}m(1-\eta^{2m-2})A_{m}-m(1-\eta^{-2(1+m)})B_{m}}\\ &\,\,\,\,\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+(1+m)(1-\eta^{2m})C_{m}-(m-1)(1-\eta^{-2m})D_{m}\bigg{]}},\end{split}$ (65)
@@
Figure 5: (a) Color map of the total growth rate $\Lambda_{2}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=0.005$ and Ca =100=100= 100. Linearly stable (S) and unstable (U) regions are shown in blue and red, respectively, and the green curve shows the marginal stability curve. Effect of (b) $m$ and (c) Ca on the marginal stability curve in (a). Panels (d), (e), and (f) show the same as (a), (b), and (c) for a contractile system ($S=-1$).
@@
The dependence of $\Xi^{c}_{m}$, $\Xi^{a}_{m}$ and $\Xi^{e}_{m}$ on $m$, $\eta$ and $\lambda$ is illustrated in Fig. 4, where panels (a), (b) and (c) show the dependence on drop size $\eta$ for distinct modes $m$, and panels (d), (e) and (f) show the effect of viscosity ratio $\lambda$ for mode $m=2$. Capillary stresses always favor a circular drop and are therefore stabilizing, especially in the case of small drops and high wavenumbers; this damping effect of surface tension is most pronounced in low-viscosity drops. In the case of extensile systems, activity is destabilizing in the case of drops of small to intermediate sizes ($0\leq\eta\lesssim 0.5$), and stabilizing for large drops, and the range of unstable sizes increases with $m$; these trends are reversed for contractile systems. In both cases, low-viscosity drops are the ones that are most affected by activity (largest magnitude of the growth rate), and the limit of high-viscosity drops ($\lambda\to\infty$) is neutrally stable. On the other hand, nematic elasticity is always stabilizing, especially in the case of small drops. These effects are most pronounced at low wavenumbers and low viscosity ratios.
@@
Figure 5(a) shows a color plot of the total growth rate $\Lambda_{2}$ corresponding to mode $m=2$ in the $(\eta,\lambda)$ parameter space, for an extensile system. The system is stable at lower values of the radius ratio ($0\leq\eta\lesssim 0.1$) followed by a range of unstable droplet sizes $(0.1\lesssim\eta\lesssim 0.25)$, and eventually regains stability at larger values of $\eta$. This behavior can be explained as follows: at lower values of $\eta$, capillary and elastic stresses dominate and are stabilizing; as $\eta$ increases, their effect weakens and the system becomes unstable as a result of active stresses; at larger values of $\eta$, active stresses become stabilizing leading to a negative growth rate. The marginal stability curve demarcating these regimes is only weakly affected by the viscosity ratio. As shown in Fig. 5(b), increasing the mode number $m$ tends to increase the size of the unstable region. Expectedly, decreasing the capillary number tends to stabilize the system across all drop sizes, especially at low viscosity ratios as illustrated in Fig. 5(c). The contractile case is analyzed similarly in Fig. 5(d,e,f). In this case, alternating regions of stable and unstable growth rates are not observed; rather, a single stable region at low values of $\eta$ gives way to an unstable region in the case of larger drops.
@@
$f^{e}_{rm}=\frac{m^{2}\eta^{m-1}}{\eta^{2m}-1}\bigg{(}(2-m)r^{m-3}-\frac{m+2}{r^{m+3}}\bigg{)}\cos m\theta,\qquad f^{e}_{\theta m}=\frac{m^{3}\eta^{m-1}}{\eta^{2m}-1}\bigg{(}r^{m-3}-\frac{1}{r^{m+3}}\bigg{)}\sin m\theta.$ (67)
They are plotted in Fig. 6 for mode $m=5$ and $\eta=0.45$, along with the first-order surface tractions acting on the interface. The fields shown in the figure suggest a destabilizing effect of bulk stresses and a stabilizing effect of surface tractions, both elastic and active (in the extensile case). The interplay of these forces, together with hydrodynamic stresses whose distribution is not explicitly known, gives rise to the trends found in Figs. 4 and 5.
Figure 6: (a) Active and (b) elastic bulk force densities and surface tractions at linear order in $\epsilon$ for the deformation mode $m=5$, in the case of a drop with radius $\eta=0.45$. In each case, the left plot shows the first-order force density, $\mathbf{f}^{a,e}_{1}=\nabla\cdot\bm{\sigma}^{a,e}_{1}$, in the bulk of the nematic, while the right plot shows the first-order traction field, $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a}$ or $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{e}+\mathbf{t}_{s,1}^{e}$, on the drop surface. Panel (a) shows the active force and traction for an extensile system with $S=+1$.
@@ IV.1 Governing equations, boundary conditions and non-dimensionalization
$\mathbf{n}\cdot(\bm{\sigma}^{h}-\overline{\bm{\sigma}}^{a}-\overline{\bm{\sigma}}^{e}-\overline{\bm{\sigma}}^{h}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,+\,\overline{\mathbf{t}}_{s}^{e}}=\frac{1}{\mathrm{Ca}}\kappa\mathbf{n}.$ (68)
Given our choice of the normal as pointing outward, the elastic traction exerted by the nematic drop on the interface is now given by
$\overline{\mathbf{t}}_{s}^{e}=-\xi^{2}\frac{\partial}{\partial s}\bigg{(}\frac{\partial\beta}{\partial n}\mathbf{n}\bigg{)},$ (69)
where arclength is still defined in the counterclockwise direction.
@@ IV.2.1 Base state
In the base state, the nematic field lines inside the drop are also circles centered at the origin, with $\beta_{0}$ given by Eq. (19). A topological defect of charge +1+1+ 1 is present at the origin. The base state active and elastic stresses are still given by Eq. (20). They induce a pressure field inside the drop that diverges at the location of the defect, while the pressure outside is uniform:
$\overline{p}_{0}(r)=S\ln r-\frac{\xi^{2}}{2r^{2}},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{0}=S\ln\eta-\frac{\xi^{2}}{2\eta^{2}}-\frac{1}{\eta\mathrm{Ca}}}.$ (70)
@@ IV.2.2 Domain perturbation and linearized problem
$\beta_{1}(r,\theta)=\widetilde{\zeta}_{1}m\eta^{-m-1}r^{m}\sin m\theta.$ (71)
@@
$\begin{split}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}-\overline{\bm{\sigma}}^{a}_{1}-\overline{\bm{\sigma}}^{e}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}&=\frac{\kappa_{1}\hat{\mathbf{r}}}{\mathrm{Ca}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{\widetilde{\zeta_{1}}}{\eta}S\big{(}\cos m\theta\,\hat{\mathbf{r}}+m\sin m\theta\,\hat{\bm{\theta}}\big{)}}\\ &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\,\,\,\,\,+\widetilde{\zeta_{1}}\frac{\xi^{2}}{\eta^{3}}\Big{[}\big{(}m^{3}+m^{2}-1\big{)}\cos m\theta\,\hat{\mathbf{r}}+m^{2}\sin m\theta\,\hat{\bm{\theta}}\Big{]}\qquad\mathrm{at}\,\,\,r=\eta.}\end{split}$ (72)
@@ IV.2.3 Lorentz reciprocal theorem
Application of the Lorentz reciprocal theorem follows the same steps as discussed in Sec. III.5 and uses the same auxiliary Newtonian flow problem. Once again, the main difference with Sec. III.5 is that active and elastic stresses now appear inside the drop. As a result, the analog of Eq. (45) has no bulk contribution and reads
@@
$\begin{split}\int_{\mathcal{C}_{0}}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}-\overline{\bm{\sigma}}^{e}_{1}-\overline{\bm{\sigma}}^{a}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}\cdot\mathbf{U}\,\mathrm{d}\ell&=m\pi\Sigma_{0}\widetilde{\zeta}_{1}\bigg{\{}\frac{m^{2}-1}{\mathrm{Ca}}\frac{\overline{G}_{m}(\eta)}{\eta^{2}}+S\left(\overline{G}_{m}^{\prime}(\eta)-\frac{\overline{G}_{m}(\eta)}{\eta}\right)\\ &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\,\,\,\,\,-\frac{\xi^{2}}{\eta^{2}}\bigg{[}m\overline{G}_{m}^{\prime}(\eta)-(m^{3}+m^{2}-1)\frac{\overline{G}_{m}(\eta)}{\eta}\bigg{]}\bigg{\}}}.\end{split}$ (76)
@@
$\begin{split}\int_{\Omega_{0}^{-}}\big{(}\overline{\bm{\sigma}}^{e}_{1}+\overline{\bm{\sigma}}^{a}_{1}\big{)}:\overline{\mathbf{E}}\,\mathrm{d}S&=-\pi\Sigma_{0}\widetilde{\zeta}_{1}m{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta^{-m-1}}\left\{S\int_{0}^{\eta}r^{m}\bigg{[}r\overline{G}_{m}^{\prime\prime}(r)-\overline{G}_{m}^{\prime}(r)+m^{2}\frac{\overline{G}_{m}(r)}{r}\bigg{]}\mathrm{d}r\right.\\ &\,\,\,\,\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\xi^{2}\int_{0}^{\eta}r^{m-2}\bigg{[}mr\overline{G}_{m}^{\prime\prime}(r)+(2m^{2}-m)\overline{G}_{m}^{\prime}(r)+(m^{3}-2m^{2})\frac{\overline{G}_{m}(r)}{r}\bigg{]}\mathrm{d}r\bigg{\}}}.\end{split}$ (77)
@@ IV.3 Dispersion relation and growth rates
Inserting Eqs. (48), (76) and (77) into Eq. (75) provides an expression for the growth rate in the form on Eq. (54), with contributions from capillary, active and elastic stresses that we discuss next.
In the translational mode ($m=1$), capillary stresses do not play a role, and the active and elastic growth rates are
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi_{1}^{a}(\eta,\lambda)=3\eta\overline{A}_{1}},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi_{1}^{e}(\eta,\lambda)=-\frac{6}{\eta}\overline{A}_{1}}.$ (78)
These growth rates are plotted in Fig. 7 vs $\eta$ and $\lambda$. Active stresses are destabilizing in extensile systems and stabilizing in contractile systems, whereas elasticity always resists translation. Both active and elastic effects are most pronounced for small drops and at low viscosity ratios. In extensile systems with finite elasticity, the stability of the system in the $(\eta,\lambda)$ parameter space is shown in Fig. 7(c), where the destabilizing effect of active stresses manifests itself above a critical radius ratio whose value depends on $\xi$. Contractile systems in the presence of elasticity are always stable to translation.
@@
$\Xi_{m}^{a}(\eta,\lambda)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m\eta^{m-2}\left[2(m-1)\overline{A}_{m}+(2m+1)\eta^{2}\overline{C}_{m}\right]},$ (80)
$\Xi_{m}^{e}(\eta,\lambda)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m\eta^{m-4}\left[(m^{3}-2m^{2}-1)\overline{A}_{m}+(m^{3}-2m^{2}-4m-1)\eta^{2}\overline{C}_{m}\right]}.$ (81)
@@
Figure 7: (a) Active contribution $\Xi^{a}_{1}$ and (b) elastic contribution $\Xi^{e}_{1}$ to the growth rate $\Lambda_{1}$ for the translational mode ($m=1$), plotted as functions of $\eta$ for different viscosity ratios $\lambda$. (c) Color map of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=0.1$. Linearly stable (S) and unstable (U) regions are shown in blue and red, respectively, and the green curve shows the marginal stability curve.
The variation of the active and elastic contributions to the growth rate $\Lambda_{m}$ for deformation modes of increasing wavenumber $m$ is shown in Fig. 8(a,b). In extensile systems, activity is stabilizing in small drops and destabilizing in larger drops, especially at low wavenumbers. Elasticity, on the other hand, is always stabilizing and its effect is most pronounced in smaller drops at low wavenumbers. The effect of viscosity ratio is analyzed in Fig. 8(c,d) for mode $m=2$, which shows that low viscosity drops are the most affected by both activity and elasticity.